OpenStudy (anonymous):

When students at Luray Middle School were surveyed, 22% said they would like hamburgers at lunch rather than hotdogs. What is that percent rounded to the nearest compatible fraction? A) 2/5 B) 1/22 C) 1/5 D) 1/2

OpenStudy (anonymous):

C, b/c 22%= 22/100 approximately equals to 2/10= 2/5

OpenStudy (anonymous):

*1/5

OpenStudy (anonymous):

Thanks @sumthenoob

OpenStudy (anonymous):

youre wlc :p